# MOCHa: Multi-Objective Reinforcement Mitigating Caption Hallucinations

### v1

@@ C.2 Additional Ablations
KL Penalty Ablation. We demonstrate the effect of our KL penalty in the reward function by performing MOCHa optimization without this term. As can be observed in the fifth row of Table 5, optimization without this penalty improves the NLI-based reward $\bar{p}$ while degrading other measures of text quality (including non-optimized metrics like CIDEr and SPICE). We hypothesize that allowing the model to freely deviate from its initial distribution encourages it towards a degenerate solution with respect to $\bar{p}$, which may be the easiest reward term to over-optimize in an unconstrained setting. This is also reflected qualitatively as seen in Figure 11. As illustrated in the figure, captions generated by the model trained without the KL penalty ($-r_{kl}$) do not contradict the image, but rather contain generic text (e.g. a painting with some items), lacking adequate detail. By contrast, optimizing with the KL penalty reward yields captions that are both descriptive and consistent with the input condition, reflected in the improved scores across metrics in Table 5 and the quality of predictions of the full reward model ($r$) in Figure 11. This is attributed to the ability of the KL penalty to mitigate over-optimization, which benefits both optimized rewards.
@@
PPO Ablation. We also ablated the selection of RL algorithm, by replacing PPO with the SCST algorithm upon which it is based (noting that SCST is the common name for the REINFORCE algorithm in the context of image captioning) [40, 33, 30]. As is seen in Table 5, PPO outperforms SCST across metrics, consistent with prior work on PPO finding that it avoids instabilities during optimization that may allow it to converge to a more optimal solution [33, 26, 49].
C.3 Additional Comparisons
@@
Comparison to VLMs Finetuned with RLHF. LLaVa-RLHF [38] is a concurrent work, which aims to reduce hallucinations in instruction tuned models using factually-grounded RLHF. In Table 6, we provide a quantitative comparison between LLaVa-RLHF and BLIP+MOCHa over 100 samples of the OPENChair dataset. For LLaVa-RLHF decoding we use both stochastic sampling with the default parameters recommended by the authors, as well as greedy sampling (as beam search is not implemented for LLaVa-RLHF). For a fair comparison, we use greedy decoding for BLIP+MOCHa as well. As LLaVa-RLHF tends to generate long paragraphs which follow an image description with subjective commentary, we terminate generation after a single sentence, which usually corresponds to an image caption. The instruction given to LLaVa-RLHF is “describe the image briefly”. As seen in the table, our method outperforms LLaVa-RLHF by this measure of open-vocabulary hallucinations. This is further seen in Figure 12, which shows example captioning predictions for these models, illustrating that LLaVa-RLHF may be more prone to hallucinations.

### v3

@@ C.3 Additional Ablations
We demonstrate the effect of our KL penalty in the reward function by performing MOCHa optimization without this term. As can be observed in the fifth row of Table 7, optimization without this penalty improves the NLI-based reward $\bar{p}$ while degrading other measures of text quality (including non-optimized metrics like CIDEr). We hypothesize that allowing the model to freely deviate from its initial distribution encourages it towards a degenerate solution with respect to $\bar{p}$, which may be the easiest reward term to over-optimize in an unconstrained setting. This is also reflected qualitatively as seen in Figure 12. As illustrated in the figure, captions generated by the model trained without the KL penalty ($-r_{kl}$) do not contradict the image, but rather contain generic text (e.g. a painting with some items), lacking adequate detail. By contrast, optimizing with the KL penalty reward yields captions that are both descriptive and consistent with the input condition, reflected in the improved scores across metrics in Table 7 and the quality of predictions of the full reward model ($r$) in Figure 12. This is attributed to the ability of the KL penalty to mitigate over-optimization, which benefits both optimized rewards.
@@
PPO Ablation. We also ablated the selection of RL algorithm, by replacing PPO with the SCST algorithm upon which it is based (noting that SCST is the common name for the REINFORCE algorithm in the context of image captioning) Sutton and Barto (2018); Schulman et al. (2017); Rennie et al. (2017). As is seen in Table 7, PPO outperforms SCST across metrics, consistent with prior work on PPO finding that it avoids instabilities during optimization that may allow it to converge to a more optimal solution Schulman et al. (2017); Ouyang et al. (2022); Ziegler et al. (2020).
C.4 Additional Comparisons
Comparison to Dedicated Models In Table 8 we provide full numeric results for older dedicated models compared to a modern VLM without further optimization, showing that they are outperformed by all metrics.
@@ C.4 Additional Comparisons
Comparison to RLHF-Tuned VLMs. LLaVa-RLHF Sun et al. (2023a) is a concurrent work, which aims to reduce hallucinations in instruction tuned models using factually-grounded RLHF. In Table 10, we provide a quantitative comparison between LLaVa-RLHF and BLIP+MOCHa over 100 samples of the OPENChair benchmark. For LLaVa-RLHF decoding we use both stochastic sampling with the default parameters recommended by the authors, as well as greedy sampling (as beam search is not implemented for LLaVa-RLHF). For a fair comparison, we use greedy decoding for BLIP+MOCHa as well. As LLaVa-RLHF tends to generate long paragraphs which follow an image description with subjective commentary, we terminate generation after a single sentence, which usually corresponds to an image caption. The instruction given to LLaVa-RLHF is “describe the image briefly". As seen in the table, our method outperforms LLaVa-RLHF by this measure of open-vocabulary hallucinations. This is further seen in Figure 13, which shows example captioning predictions for these models, illustrating that LLaVa-RLHF may be more prone to hallucinations.
@@
MOCHa’s Improvement (OCH) in %
Model     without filtering with filtering
BLIP-B 4.9% 4.8%
BLIP-L 2.0% 2.3%
BLIP2 7.3% 6.9%
GIT-B 7.0% 7.1%
Table 12: Performance of MOCHa with and without manual filtering. We compare performance on the OpenCHAIR (OCH) benchmark before and after it is manually filtered, as measured by the improvement provided by MOCHa on OpenCHAIR scores across various models. We observe similar results before and after filtering, corresponding to the relative high quality of the generated data and consistent with the small proportion of data that was removed.
Appendix D Extended Discussion of Previous Work
We provide here an extended discussion of related methods, shown in Figure 10.
D.1 Similarity Based Metrics
CLIPScore Hessel et al. (2022) propose CLIP cross-modal similarity for detecting mismatches between text and images, including hallucinations, and Shi et al. (2022) propose a similar embedding-based metric for video captioning. However, Xu et al. (2023) find that CLIP tends to assign high similarity to texts with minor modifications (“hard negatives”) that contradict the corresponding image. The Egoshots Semantic Fidelity metric Agarwal et al. (2020) and VIFIDEL Madhyastha et al. (2019) use embedding similarity between object annotations or detections in images and items in predicted captions. FAIEr Wang et al. (2021) proposes a learned fidelity metric, which must be trained on automatically-generated scene graphs. Unlike these methods, our benchmark provides an explicit measure of hallucinations that can be directly examined (predicted captions on the OpenCHAIR benchmark images).
D.2 Closed Vocabulary Algorithms
UD-L Biten et al. (2021) identifies object hallucinations with bias towards the prior distribution of objects in context found in the training data, and proposes the use of synthetically debiased captions. CIIC Liu et al. (2022) focuses on captioning models with a closed-vocabulary object detection backbone, inserting components into the object detector and text decoder to reduce spurious correlations. TLC Petryk et al. (2023) proposes a text decoding method applied to existing captioning models, to avoid generating COCO object tokens if they have insufficient confidence. The more recent work ObjMLM Dai et al. (2023) proposes masking objects from closed vocabulary lists as a training objective. The concurrent work Woodpecker Yin et al. (2023) combines closed-vocabulary object detection with LLM-guided decoding to avoid hallucinations in generated text. Unlike these works, our MOCHa optimization method does not rely on a closed list of object types.